OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: 9x − 5y = 12 2y + 4 = 6x (-2,-6) (2/9, -2) (3, 5) (-1/3, -3)

OpenStudy (anonymous):

Hey, do you have a graphing caluclator?

OpenStudy (anonymous):

no,:(

OpenStudy (anonymous):

Ok, go on google and type in free graphing calculator. Graph these two equations and find the point that they intersect. This point is your solution.

OpenStudy (anonymous):

-1/3,-3

OpenStudy (anonymous):

Do you know how to solve simultaneous equations? Either use substitution method or elimination.

OpenStudy (anonymous):

okay thanks guys(:

OpenStudy (anonymous):

Using substitution method, \[9x-5y=12\] \[6x-2y=4 \] So \[2y= 4-2x \]\[y= 2-x\] \[9x-5(2-x)= 12 \]Solve for x and then y